OpenStudy (anonymous):

For a certain drug, the rate of reaction in appropriates units is given by R'=7/t + 2/t^2, when t is measured in hours after the drug is administed. Find the total reaction to the drug from t=5 to t=10

OpenStudy (anonymous):

integrate it from 5 to 10

OpenStudy (anonymous):

\[\int\limits_{10}^{5} 7\ln \left| x \right| + 2/t\]

OpenStudy (anonymous):

?

OpenStudy (amistre64):

\[\int\limits_{5}^{10} \frac{7}{t} + \frac{2}{t^2} dt \rightarrow \rightarrow 7\ln(t) - \frac{2}{t}\]

OpenStudy (amistre64):

1/t^2 integrates to (-1/t) :) keep track of those signs

OpenStudy (anonymous):

yeah those signs are gonna kill lol

OpenStudy (amistre64):

15.918 - 10.866 = 5.052 give or take